OpenStudy (anonymous):

What are the possible numbers of positive, negative, and complex zeros of f(x) = -3x^4 + 5x^3 - x^2 + 8x + 4?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

fractions where the numerator divides 4 and the denominator divides 3

OpenStudy (anonymous):

there are a few of those

OpenStudy (anonymous):

the 8x and 4?

OpenStudy (anonymous):

divisors of 4 are \[\pm1,\pm2,\pm4\] so do the same with a 3 in the denominator

OpenStudy (anonymous):

ignore everything but the 4 at the end and the 3 at the beginning

OpenStudy (anonymous):

k so since you already found the 4 all i gotta do is the -3x^4?

OpenStudy (anonymous):

the others are \[\pm\frac{1}{3}\] and so on

OpenStudy (anonymous):

do you want to see the answer choices?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i don't know how to do this im not sure what you mean by 3 as the denominator like do you want any fraction with 3 as the denominator ?

OpenStudy (anonymous):

\[\pm \frac{1}{3},\pm \frac{2}{3},\pm\frac{4}{3}\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yeah so like [\pm\frac{ 6 }{ 3 }\] would that work?

OpenStudy (anonymous):

\[\pm \frac{ 6 }{ 3 }\]

OpenStudy (anonymous):

lol just the ones i wrote the top as to divide 4

OpenStudy (anonymous):

okay so i devide all those by 4 or what?

OpenStudy (anonymous):

let me write them out for you ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\pm1,\pm2,\pm4\pm \frac{1}{3},\pm \frac{2}{3},\pm\frac{4}{3}\]

OpenStudy (anonymous):

top goes in to 4, bottom goes in to 3

OpenStudy (anonymous):

ohhh i see what you mean like the numerator goes into 4 and he denominator goes into 3?

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

but thats not one of the answer choices

OpenStudy (anonymous):

oooh i see

OpenStudy (anonymous):

crap we have to use descartes rule of sign

OpenStudy (anonymous):

\[ f(x) = -3x^4 + 5x^3 - x^2 + 8x + 4\] has 3 changes in sign from -3 to +5, from +5 to -1, from -1 to +8

OpenStudy (anonymous):

so there are either 3 positive zeros, or one positive zero (you count down by twos)

OpenStudy (anonymous):

\[f(-x) = -3x^4 - 5x^3 - x^2 - 8x + 4\] has one change in sign from -8 to +4

OpenStudy (anonymous):

i thought 0 was a nutreul number? not negative nor positive?

OpenStudy (anonymous):

changes in sign of the coefficients yes, zero is not negative or positive, we are counting the roots

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Positive: 2 or 0; negative: 2 or 0; complex: 4 or 2 or 0 Positive: 1; negative: 3 or 1; complex: 2 or 0 Positive: 3 or 1; negative: 1; complex: 2 or 0 <--- s this the answeR? Positive: 4 or 2 or 0; negative: 2 or 0; complex: 4 or 2 or 0

OpenStudy (anonymous):

well we don't have to do much work because only one choice starts with positive zeros 3 or 1

OpenStudy (anonymous):

yeah C as you picked

OpenStudy (anonymous):

thank you i got a whole bunch more btw Xd